OpenStudy (anonymous):

Explain why x = 3 makes 4x − 1 ≤ 11 true but not 4x − 1 < 11?

OpenStudy (igreen):

Plug it in. \(4x - 1 \le 11\) \(4(3) - 1 \le 11\)

OpenStudy (anonymous):

? @iGreen

OpenStudy (igreen):

Multiply 4 * 3 and subtract 1

OpenStudy (anonymous):

11 @iGreen

OpenStudy (igreen):

Yes, that gives us: \(11 \le 11\)

OpenStudy (haseeb96):

so 11 will be equal or greater from 11

OpenStudy (igreen):

And for the other equation it would give us: 11 < 11

OpenStudy (haseeb96):

?@sjman1

OpenStudy (igreen):

Why is \(11 \le 11\) true and why is \(11 < 11\) not true?

OpenStudy (anonymous):

do i put that down as my awnswer? @iGreen

OpenStudy (igreen):

Yeah, say \(4x - 1 \le 11\) equals \(11 \le 11\) and \(4x - 1 < 11\) equals \(11 < 11\) So you could say the reason is because the first one has a less than OR EQUAL TO sign, and 11 is equal to 11, but the other one only has a less than sign and 11 isn't less than 11.

OpenStudy (anonymous):

thanks! :) @igr